length of rectangular garden is 6 feet less than 3 times its width the area is 360 square feet length garden is?
Answer by checkley77(12844)

L=3W-6
LW=360
(3W-6)W=360
3W^2-6W-360=0
3(W^2-2W-120)=0
3(W-12)(W+10)=0
W-12=0
W=12 ANS. FOR THE WIDTH.
L=3*12-6=36-6=30 ANS. FOR THE LENGTH.
PROOF:
30*12=360
360=360